# A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

### v1

@@ 1.2 Tools: exact penalization, smoothing, and variance reduction
A key tool in our approach is the exact penalization. This tool transforms problem Eq. P into an unconstrained one. Notably, when the penalty parameter is sufficiently large, the stationarity point of the penalized problem can align with that of the original constrained problem. Building on the exact penalization, we then design a stochastic FOM to solve the resulting unconstrained problem.
Using the penalty function $h=[\cdot]_{+}:a\to\max\{a,0\}$ for an inequality constraint, an exact penalty reformulation [8, 33, 38, 45] of problem Eq. P is given by
(5) $\min_{{\mathbf{x}}}\,\,F({\mathbf{x}})=f({\mathbf{x}})+\beta g_{+}({\mathbf{x}}),\text{ with }g_{+}({\mathbf{x}})=h(g({\mathbf{x}})),$
@@
where $\beta>0$ is the penalty parameter. The exact penalization results regarding global minimizers are established in [33]. Note that $\mathrm{Sign}([g(\mathbf{x})]_{+})\boldsymbol{\zeta}_{g}$ belongs to $\partial g_{+}(\mathbf{x})$, where $\boldsymbol{\zeta}_{g}\in\partial g(\mathbf{x})$. Here, $\mathrm{Sign}:\mathbb{R}\mapsto\{-1,0,1\}$ refers to the sign operator, and $\partial$ denotes the (Clarke) subdifferential [12, Section 1.2] of a continuous function. Due to the stochastic nature of $g$, it is impractical to compute the exact value of $\mathrm{Sign}([g(\mathbf{x})]_{+})\boldsymbol{\zeta}_{g}$. Approximating this value is also challenging, because $\mathrm{Sign}$ is discontinuous, making it impossible for the discrepancies between $\mathrm{Sign}([g(\mathbf{x})]_{+})$ and $\mathrm{Sign}([g(\mathbf{x},\mathcal{B})]_{+})$ to be consistently small enough. Here, $\mathcal{B}$ is a sample set of $\xi_{g}$ over the distribution $\mathbb{P}_{g}$, and
(6) $g(\mathbf{x},\mathcal{B}):=\frac{1}{|\mathcal{B}|}\sum_{\xi_{g}\in\mathcal{B}}g({\mathbf{x}},\xi_{g}),$
@@
with $|\mathcal{B}|$ denoting the cardinality of $\mathcal{B}$. Consequently, solving problem Eq. 5 directly by a stochastic method is challenging. These difficulties extend to algorithm design and convergence analysis, complicating the establishment of theoretical guarantees.
To resolve this issue, we introduce a smoothed function of $h$ with a smoothing parameter $\nu>0$ as follows,
@@
$\vspace{-3mm}h^{\nu}(z)$ $:=\max_{0\leq y\leq 1}\langle y,z\rangle-\frac{\nu}{2}y^{2}=\left\{\begin{aligned} \frac{z^{2}}{2\nu},\quad&\text{ if }0\leq z\leq\nu,\\ z-\frac{\nu}{2},&\text{ if }z>\nu,\\ 0,\quad\,\,\,&\text{ otherwise.}\end{aligned}\right.$
@@
We then present a new penalty model of problem Eq. P:
(PP) $\min_{{\mathbf{x}}}\,\,F^{\nu}({\mathbf{x}})=f({\mathbf{x}})+\beta h^{\nu}(g({\mathbf{x}})).$
@@
Since $h^{\nu}$ is smooth and $\nabla h^{\nu}(z)=\mathrm{Proj}_{[0,1]}\left(\frac{z}{\nu}\right)$, we can access a subgradient of $h^{\nu}(g({\mathbf{x}}))$ via $\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g}$, where $\boldsymbol{\zeta}_{g}\in\partial g(\mathbf{x})$. In addition, as $\mathrm{Proj}_{[0,1]}$ is Lipschitz continuous, the difference between $\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g}$ and $\mathrm{Proj}_{[0,1]}\left(\frac{u}{\nu}\right)\boldsymbol{\zeta}_{g}$ can be well controlled if $u$ is a good approximation of $g({\mathbf{x}})$.
@@
We apply a SPIDER-type (or SARAH) [16, 32] variance reduction technique to estimate $g({\mathbf{x}}^{(k)})$ to obtain $u^{(k)}$ and perform the update in Eq. 1 for each $k$. The variance reduction tool, together with the smoothing tool, enables us to achieve better complexity results for our method than state-of-the-art results to produce an $(\epsilon,\epsilon)$-KKT point of (P).
@@ 1.3 Contributions
First, to address the expectation constraint involved in Eq. P, we introduce a new smooth penalty function $h^{\nu}\circ g$. Leveraging this function, we propose solving problem Eq. PP. For any given $\epsilon>0$, we demonstrate that, under easily verifiable conditions, any near-$\epsilon$ stationary point of problem Eq. PP (in expectation) is an $(O(\epsilon),O(\epsilon))$-KKT point of problem Eq. P (in expectation). This property enables us to pursue an $(\epsilon,\epsilon)$-KKT point of problem Eq. P by finding a near-$O(\epsilon)$ stationary point of problem Eq. PP. Our framework is well-suited for many applications such as problem Eq. 4, which involves a nonconvex objective and convex expectation constraints satisfying the required assumptions.
Second, we exploit the structure of problem Eq. PP and propose a single-loop SPIDER-type stochastic FOM to solve it, with updates given in Eq. 1. At each iteration, the SPIDER technique [16] is employed to generate an approximation $u^{(k)}$ of the constraint function value. Using the tool of Moreau envelope, we demonstrate that the proposed method can, in expectation, find a near-$\epsilon$ stationary point of problem Eq. PP within $O(\epsilon^{-4})$ iterations. This result matches the known lower bound for such tasks [19], highlighting the efficiency of our approach. Combining this with the theory of exact penalization, we further analyze the complexity of obtaining an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation.
Based on the structure of $g$, we categorize the complexity results of our method as follows:
@@
    A generic stochastic $g$: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\epsilon^{-6})$
@@
    A finite-sum structured [^1] Here, we assume that $\mathbb{P}_{g}$ is a uniform distribution over a finite set of size $N$, where $N<\epsilon^{-4}$. $g$: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\sqrt{N}\epsilon^{-4})$.
When $g$ is generic stochastic, our complexity result is lower than state-of-the-art results by a factor of $\epsilon^{-2}$ in either OGC or CFC; when $g$ enjoys a finite-sum structure, the highest order among our results of OGC, CGC, and CFC is $O(\sqrt{N}\epsilon^{-4})$, representing an improvement over the previously best-known $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$ bound. Detailed comparisons are provided in Table 1. In this table, the sharpness assumption and the gradient-away assumption that we make are weaker than the (uniform) Slater-type CQ in [18, 30], which is weaker than a strong feasibility condition (see Eqn. Eq. 41 below) in [9]. These assertions will be verified in Appendix A.
Third, the ability to handle nonsmooth problems using a single-loop approach eliminates the need for tuning inner iterations, distinguishing the proposed method from double-loop methods, such as those in [9, 24, 30]. This simplification reduces computational overhead and makes the method more accessible for real-world applications. In numerical experiments, we test both the deterministic and stochastic versions of our method, for two problems across two datasets. Notably, the deterministic version is significantly faster than state-of-the-art methods, and the stochastic version achieves even more impressive speedups.
Methods SL structure of $g$ CQ on $g$ OGC CGC CFC
IPP [9] no stochastic convex $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
finite-sum convex MFCQ, and $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
stochastic WC strong feasibility $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
finite-sum WC $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
SSG [18] yes stochastic convex Slater-type CQ $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-8})$
@@
Ours yes stochastic convex sharpness, or $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
finite-sum convex gradient away $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
stochastic WC gradient away, and $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
finite-sum WC $B_{g}>\rho_{g}D^{2}$ $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
finite-sum WC Slater-type CQ $O(\epsilon^{-4})$ $O(N\epsilon^{-4})$ $O(N\epsilon^{-4})$

### v2

@@ 1.2 Tools: exact penalization, smoothing, and variance reduction
A key tool in our approach is the exact penalization. This tool transforms problem Eq. P into a problem without function constraints. Notably, when the penalty parameter is sufficiently large, the stationary point of the penalized problem can align with the KKT point of the original problem. Building on the exact penalization, we then design a stochastic FOM to solve the resulting problem.
Let the penalty function $h({\mathbf{z}})={\mathbf{e}}^{\top}[{\mathbf{z}}]_{+}$, where ${\mathbf{e}}\in\mathbb{R}^{m}$ is the all-ones vector and ${\mathbf{z}}]_{+}:=(\max\{z_{1},0\},\ldots,\max\{z_{m},0\})^{\top}$ for any ${\mathbf{z}}\in\mathbb{R}^{m}$. Then an exact penalty reformulation [9, 50, 51, 57, 62, 65] of problem Eq. P is
(6) $\min_{{\mathbf{x}}\in\mathcal{X}}\,\,F({\mathbf{x}}):=f({\mathbf{x}})+\beta g_{+}({\mathbf{x}}),\text{ with }g_{+}({\mathbf{x}})=h({\mathbf{g}}({\mathbf{x}}))=\sum_{i=1}^{m}[g_{i}({\mathbf{x}})]_{+},$
@@
where $\beta>0$ is the penalty parameter. The exact penalization results regarding global minimizers are established in [50]. Note that $\sum_{i=1}^{m}\mathrm{Sign}([g_{i}(\mathbf{x})]_{+})\boldsymbol{\zeta}_{g_{i}}$ belongs to $\partial g_{+}(\mathbf{x})$, where $\boldsymbol{\zeta}_{g_{i}}\in\partial g_{i}(\mathbf{x})$, and $\mathrm{Sign}:\mathbb{R}\mapsto\{-1,0,1\}$ refers to the sign operator. Due to the stochastic nature of $\mathbf{g}$}, it is impractical to compute the exact value of $\mathrm{Sign}([g_{i}(\mathbf{x})]_{+})$ for each $i$. Approximating this value is also challenging, because $\mathrm{Sign}$ is discontinuous, making it impossible for the discrepancies between $\mathrm{Sign}([g_{i}(\mathbf{x})]_{+})$ and $\mathrm{Sign}([g_{i}(\mathbf{x},\mathcal{B})]_{+})$ to be consistently small enough. Here, $\mathcal{B}$ is a sample set of $\xi_{{\mathbf{g}}}$ over the distribution $\mathbb{P}_{{\mathbf{g}}}$, and we define
(7) $g_{i}(\mathbf{x},\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{\xi_{{\mathbf{g}}}\in\mathcal{B}}g_{i}({\mathbf{x}},\xi_{{\mathbf{g}}}),\forall\,i\in[m],\quad{\mathbf{g}}({\mathbf{x}},{\mathcal{B}})=\left(g_{1}(\mathbf{x},\mathcal{B}),\ldots,g_{m}(\mathbf{x},\mathcal{B})\right)^{\top},$
@@
with $|\mathcal{B}|$ denoting the cardinality of $\mathcal{B}$. Consequently, solving problem Eq. 6 directly by a stochastic method is challenging.
To resolve this issue, we employ the Huber-type [^1] The smoothing technique is a powerful tool for handling nonsmooth problems, and we refer interested readers to [11, 12, 15] for further details and illustrative examples. Specifically, the function $\cdot\,$$_{+}$ admits various smooth approximations, including the Softplus function [36, 37], the square-root function [14], and the log-sum-exp function [15, 41]. Our complexity results remain valid as long as the smoothing function $h^{\nu}$ is Lipschitz continuous, its gradient has a Lipschitz constant that grows linearly with $1/\nu$, and the composition $h^{\nu}({\mathbf{g}}(\cdot))$ is weakly convex. The mentioned Huber-type, Softplus and log-sum-exp smoothing functions satisfy all these conditions. We choose the Huber-type smoothing because its gradient takes the particularly simple form $\mathrm{Proj}_{[0,1]}(\cdot/\nu)$, which facilitates a clean and efficient algorithmic implementation. smoothing [16, 48, 55] function $h^{\nu}$ given as follows,
@@
$\vskip-8.53581pth^{\nu}({\mathbf{z}})$ $:=\sum_{i=1}^{m}H^{\nu}(z_{i}),\text{ with }H^{\nu}(z_{i})=\max_{0\leq y\leq 1}\left[\langle y,z_{i}\rangle-\frac{\nu}{2}y^{2}\right]=\left\{\begin{aligned} \frac{z_{i}^{2}}{2\nu},\quad&\text{ if }0\leq z_{i}\leq\nu,\\ z_{i}-\frac{\nu}{2},&\text{ if }z_{i}>\nu,\\ 0,\quad\,\,\,&\text{ otherwise},\end{aligned}\right$.
@@
where $\nu>0$ is a smoothing parameter. Then instead of directly working on Eq. 6, we present a new penalty model of problem Eq. P:
($P$) $\min_{{\mathbf{x}}\in\mathcal{X}}\,\,F^{\nu}({\mathbf{x}})=f({\mathbf{x}})+\beta h^{\nu}({\mathbf{g}}({\mathbf{x}})).$
@@
Since $h^{\nu}$ is smooth and $\nabla h^{\nu}({\mathbf{z}})=\mathrm{Proj}_{[0,1]^{m}}\left(\frac{{\mathbf{z}}}{\nu}\right)$, we can access a subgradient of $h^{\nu}({\mathbf{g}}({\mathbf{x}}))$ via $\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}$, where $\boldsymbol{\zeta}_{g_{i}}\in\partial g_{i}(\mathbf{x})$ for each $i\in[m]$. In addition, as $\mathrm{Proj}_{[0,1]}$ is Lipschitz continuous, the difference between $\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}$ and $\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}$ can be well controlled if $u_{i}$ is a good approximation of $g_{i}({\mathbf{x}})$.
@@
For each $i\in[m]$ and each $k$, we apply a SPIDER-type (or SARAH) [24, 49] variance reduction technique to estimate $g_{i}({\mathbf{x}}^{(k)})$ to obtain $u_{i}^{(k)}$ and perform the update in Eq. 2. The variance reduction tool, together with the smoothing tool, enables us to achieve better complexity results for our method than state-of-the-art results to produce an $(\epsilon,\epsilon)$-KKT point of Eq. P.
@@ 1.3 Contributions
First, to address the expectation constraint involved in Eq. P, we introduce a Huber-type smoothing penalty function $h^{\nu}\circ{\mathbf{g}}$. Leveraging this function, we propose solving problem Eq. PP. For any given $\epsilon>0$, we demonstrate that, under easily verifiable conditions, any near-$\epsilon$ stationary point of problem Eq. PP (in expectation) is an $(O(\epsilon),O(\epsilon))$-KKT point of problem Eq. P (in expectation). This property enables us to pursue an $(\epsilon,\epsilon)$-KKT point of problem Eq. P by finding a near-$O(\epsilon)$ stationary point of problem Eq. PP. Our framework is well-suited for many applications such as problem Eq. 5, which involves a nonconvex objective and convex expectation constraints satisfying the required assumptions.
Second, we exploit the structure of problem Eq. PP and propose a SPIDER-type stochastic FOM to solve it, with updates given in Eq. 2. At each iteration, the SPIDER technique [24] is employed to generate an approximation $u_{i}^{(k)}$ of the constraint function value $g_{i}({\mathbf{x}}^{(k)})$ for each $i\in[m]$. Using the tool of Moreau envelope, we demonstrate that the proposed method can, in expectation, find a near-$\epsilon$ stationary point of problem Eq. PP within $O(\epsilon^{-4})$ iterations. This result matches the known lower bound for such tasks [29], highlighting the efficiency of our approach. Combining this with the theory of exact penalization, we further analyze the complexity of obtaining an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation.
Based on the structure of $\mathbf{g}$}, we categorize the complexity results of our method as follows:
@@
    A generic stochastic $\mathbf{g}$}: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\epsilon^{-6})$ .
@@
    A finite-sum structured [^2] Here, we assume that $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite set of size $N$, where $N<\epsilon^{-4}$. $\mathbf{g}$}: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\sqrt{N}\epsilon^{-4})$.
When $\mathbf{g}$} is generic stochastic, our complexity result is lower than state-of-the-art results by a factor of $\epsilon^{-2}$ in either OGC or CFC; when $\mathbf{g}$} enjoys a finite-sum structure, the highest order among our results of OGC, CGC, and CFC is $O(\sqrt{N}\epsilon^{-4})$, representing an improvement over the previously best-known $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$ bound. Detailed comparisons are provided in Table 1.
Third, our approach does not need to solve subproblems by an iterative subroutine and thus eliminates the need for tuning inner iterations, distinguishing our method from those in [10, 38, 46]. This simplification reduces computational overhead and makes the method more accessible for real-world applications. In numerical experiments, we test both the deterministic and stochastic versions of our method, for three problems across three datasets. Notably, the deterministic version is significantly faster than state-of-the-art methods, and the stochastic version achieves even more impressive speedups.
Methods RIS structure of $\mathbf{g}$} CQ on $\mathbf{g}$} OGC CGC CFC
IPP [10] yes stochastic convex MFCQ, and $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
finite-sum convex Slater’s condition $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
stochastic WC MFCQ, and $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
finite-sum WC strong feasibility $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
SSG [27] no stochastic convex Slater-type CQ $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-8})$
@@
Ours no stochastic convex Slater-type CQ $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
finite-sum convex $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
stochastic WC Slater-type CQ, with $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
finite-sum WC $B_{g}>m\rho_{g}D^{2}$ $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
Table 1: Comparison of complexity results among our method and existing ones for obtaining an $(\epsilon,\epsilon)$-KKT point of problem Eq. P with a stochastic nonsmooth weakly convex objective function $f$. RIS is short for “Require Inner Solver”, and WC is for “weakly convex”. Strong feasibility implies the Slater-type CQ with any $B_{g}>0$; see Proposition 2.1(ii). The Slater-type CQ is directly assumed in [27] when $\{g_{i}\}$ are weakly convex. Under the setting where $g_{i}$ is convex for each $i\in[m]$, Slater’s condition is equivalent to the Slater-type CQ; see Proposition 2.1(i). In the finite-sum setting, we assume that $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite set of size $N<\epsilon^{-4}$. When $\{g_{i}\}_{i\in[m]}$ are weakly convex, Huang et al. [27] provide complexity results only for the case where both $f$ and $\mathbf{g}$} are deterministic, excluding the setting where $f$ is stochastic.
1.4 Notations and definitions
We use ${\mathrm{dom}}(\cdot)$ to denote the domain of a function. $[M]$ denotes $\{1,\ldots,M\}$ for a positive integer $M$. The 0-$\infty$ indicator function of a set $\mathcal{X}$} is denoted as $\iota_{\mathcal{X}}$. The $l_{2}$-norm is written as $\|\cdot\|$. $\mathrm{co}(\cdot)$} and $\mathrm{relint}(\cdot)$ represent the closed convex hull and relative interior of a set, respectively. $\mathcal{N}_{\mathcal{X}}({\mathbf{x}})$ refers to the normal cone of a closed convex set $\mathcal{X}$ at $\mathbf{x}$}. We use $\mathrm{Proj}$, $\mathrm{Prob}$, and $\mathrm{dist}$ to denote operators of the projection, the probability, and the distance, respectively.
Definition 1.1 (subdifferential).
For a convex function $f$, its subdifferential at ${\mathbf{x}}\in{\mathrm{dom}}(f)$ is defined as [18]
$\partial f({\mathbf{x}})=\big{\{}{\boldsymbol{\xi}}:f({\mathbf{x}}^{\prime})-f({\mathbf{x}})\geq\langle{\boldsymbol{\xi}},{\mathbf{x}}^{\prime}-{\mathbf{x}}\rangle\ \text{ for all }{\mathbf{x}}^{\prime}\in{\mathrm{dom}}(f)\big{\}}.$
The subdifferential of a $\rho$-weakly convex function $f$ at ${\mathbf{x}}\in{\mathrm{dom}}(f)$ is given by
$\partial f({\mathbf{x}}):=\textstyle\partial\big{(}f(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}\big{)}({\mathbf{x}})-\rho{\mathbf{x}}.$
Definition 1.2 (Moreau Envelope [21]).
For a $\rho_{\phi}$-weakly convex function $\phi$, its Moreau envelope and proximal mapping for any $t\in(0,\frac{1}{\rho_{\phi}})$ on $\mathcal{X}$ are defined by $\phi_{t}({{\mathbf{x}}})=\min_{{\mathbf{z}}\in\mathcal{X}}\left\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\right\}$, $\operatorname{Prox}_{t\phi}({{\mathbf{x}}})=\arg\min_{{\mathbf{z}}\in\mathcal{X}}\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\}.$
We give the definition of stationarity as follows. In deterministic cases, where randomness is not involved, the expectation operator $\mathbb{E}$ should be removed from the definition below.
Definition 1.3 (stationarity).
Let $\epsilon>0$. Suppose that $\{\psi_{i}\}_{i=1}^{m}$ are all $\rho_{\psi}$-weakly convex and $\phi$ is $\rho_{\phi}$-weakly convex on $\mathcal{X}$}.
$\mathrm{(a)}$ $\mathbf{x}$$^{*}$ is called a near-$\epsilon$ stationary point of $\min_{{\mathbf{x}}\in{\mathcal{X}}}\phi({\mathbf{x}})$ in expectation, if ${\mathbf{x}}^{*}\in{\mathcal{X}}$, and there exists $0<t<\frac{1}{\rho_{\phi}}$, such that $\mathbb{E}[\|\nabla\phi_{t}({\mathbf{x}}^{*})\|^{2}]\leq~\epsilon^{2}.$
$\mathrm{(b)}$ $\mathbf{x}$$^{*}$ is called an $\epsilon$-KKT point of $\min_{{\mathbf{x}}\in\mathcal{X},\boldsymbol{\psi}({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if ${\mathbf{x}}^{*}\in\mathcal{X}$ and there exists $\boldsymbol{\lambda}_{{\mathbf{x}}^{*}}\geq 0$ such that $\mathbb{E}\left[\psi_{i}({\mathbf{x}}^{*})\right]\leq\epsilon,$ $|({\lambda}_{{\mathbf{x}}^{*}})_{i}{\psi}_{i}({\mathbf{x}}^{*})|\leq\epsilon$ for all $i\in[m]$, and $\mathbb{E}\left[\textnormal{dist}(\mathbf{0},\partial\phi({\mathbf{x}}^{*})+\sum_{i=1}^{m}(\lambda_{{\mathbf{x}}^{*}})_{i}\partial\psi_{i}({\mathbf{x}}^{*})+\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{*}))\right]\leq\epsilon.$
$\mathrm{(c)}$ $\mathbf{x}$$^{*}$ is called an $(\epsilon,\epsilon)$-KKT point of $\min_{{\mathbf{x}}\in\mathcal{X},\boldsymbol{\psi}({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if ${\mathbf{x}}^{*}\in\mathcal{X}$ and $\mathbb{E}[\|{\mathbf{x}}^{*}-\widetilde{{\mathbf{x}}}^{*}\|]\leq\epsilon$ with $\widetilde{{\mathbf{x}}}^{*}$ being an $\epsilon$-KKT point of $\min_{{\mathbf{x}}\in\mathcal{X},\boldsymbol{\psi}({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation.
1.5 Organization
The rest of this paper is organized as follows. Section 2 introduces the exact penalty models. In Section 3, we propose a SPIDER-type stochastic subgradient method and establish its iteration complexity results. Numerical results are presented in Section 4, showcasing the performance of our method. Finally, Section 5 concludes the paper.
2 Model analysis
In this section, we establish the relationships among models Eq. PEq. 6 and Eq. PP, which are presented in the following diagram for the convenience of readers. Details are given in the mentioned lemma and theorem.